a tank which has an inlet valve through which it get filled in 5 hours and outlet valve through which it becomes empty in 10 hours. when both are open how long will it take to fill the tank?
Answer by mananth(16946)

Filling or draining the complete tank is considered as 1 job
filling in 5 hours
so 1/5 of the job in 1 hour.
..
draining in 10 hours
1/10 of the job in 1 hour
..
effective filling when both are open =
1/5 - 1/10 hours
10-5/50= 5/50= 1/10
when both are open
1/10 of the job is done in 1 hour
so it will take 10 hours to fill the tank